# A Survey of Behavior Foundation Model: Next-Generation Whole-Body Control System of Humanoid Robots

### v3

@@ 4.2.2 Data Bottleneck
The data bottleneck poses a fundamental constraint in developing BFMs for humanoid robots. While datasets listed in Table III have been successfully employed to train the current BFMs, their scale remains significantly smaller than the datasets used to train LLMs or large vision models. This scarcity is exacerbated when retargeting motions to specific robotic platforms [146], where subtle morphological differences can incur severe policy performance loss. Real-world robot data is even more constrained due to hardware limitations and safety concerns. The challenge compounds when considering multimodal data requirements.
@@
KIT-ML [147] 3911 11.2 2016
AMASS [148] 11265 40.0 2019
LAFAN [149] 77 4.6 2020
BABEL [150] 13220 43.5 2021
Posescript [151] - - 2022
HumanML3D [152] 14616 28.6 2022
Motion-X [153] 81084 144.2 2023
Motion-X++ [154] 120462 180.9 2025
@@
Current BFMs predominantly rely on proprioceptive inputs, and integrating exteroceptive sensing for real-world deployment introduces new bottlenecks. Pioneering work like OmniH2O [155] explores multimodal data collection through teleoperation and pairing head-mounted RGBD cameras with whole-body proprioception and motor commands. While this enables tasks like vision-based autonomy and imitation learning, its 40-minute dataset (OmniH2O-6) remains limited to lab environments and lacks diverse contact dynamics. Currently, there is no existing dataset that provides large-scale and temporally aligned recordings of proprioception, vision, and contact dynamics across diverse environments. Therefore, larger-scale, high-quality, and more curated datasets are urgently needed to enhance the effectiveness of current BFMs and to aid in the development of future models.
4.2.3 Embodiment Generalization
@@
Despite the use of the term ”BFM”, current BFMs are typically trained on specific humanoid robot embodiments with fixed morphology, actuator dynamics, and sensor configurations. While they excel at controlling the robot they were trained on, it is currently intractable for them to generalize to novel embodiments with different body shapes, actuator types, or degrees of freedom. This limitation stems from several key challenges in embodiment generalization. One major issue is morphological mismatch, where policies designed for a specific kinematic or dynamic structure struggle to adapt to robots with differing link lengths, joint types, or mass distributions. Additionally, differences in actuator dynamics—such as variations in torque, latency, or control modes (e.g., position vs. torque control)—can destabilize policies when transferred between systems. Sensor diversity adds another layer of complexity, as BFMs often assume consistent sensor inputs and have difficulty managing situations with missing or additional sensing modalities. Furthermore, reward functions created for one type of embodiment (e.g., humanoid walking) may not effectively translate to others (e.g., hexapod gait patterns), leading to confusion in task adaptation. These challenges underscore the necessity for more adaptable BFM architectures that can abstract skills across various robotic platforms.
5 Opportunities and Risks
@@
In this section, we explore key research opportunities that could enhance the broader adoption of BFMs, including their integration into high-level ML systems, scaling laws, post-training techniques, and multi-agent coordination. We also critically examine the risks associated with BFMs, such as ethical concerns in human-robot interaction and safety mechanisms for real-world deployment. By addressing these opportunities and risks, the robotics community can ensure that BFMs evolve into robust, adaptable, and socially responsible controllers for next-generation humanoid systems.
@@
5.1 Opportunities
5.1.1 Multimodal BFMs
@@
A promising direction for BFMs lies in expanding beyond proprioceptive inputs to multimodal sensory integration, incorporating exteroceptive signals such as vision, acoustic signals, and tactile feedback [156, 157, 158]. While current BFMs primarily rely on proprioceptive information, richer perceptual inputs could enable more robust and adaptive behaviors in unstructured environments. For example, integrating real-time visual perception could enable humanoids to dynamically adjust their movements based on object positions, terrain conditions, or human interactions, enhancing both safety and task performance. However, achieving multimodal BFMs requires the advancement of both large-scale datasets and scalable training paradigms, which are key challenges for the field.
5.1.2 High-level ML System
@@
Recent advances in foundation models like LLMs and VLAs have demonstrated their potential as controllers for coordinating specialized AI systems [159, 160, 161, 34, 162, 163, 164]. For example, HuggingGPT [159] leverages LLMs to orchestrate task planning across diverse models, while Eureka [34] employs LLMs to automate reward function design in RL. These projects present a natural pathway for integrating BFMs into high-level ML systems, where they could serve as universal low-level controllers for humanoid robots. By combining the reasoning capabilities of LLMs (for task planning and adaptation) with the motor priors encoded in BFMs (for real-time execution), such systems could achieve unprecedented flexibility in handling complex and multi-step physical tasks, while minimizing task-specific engineering. The ultimate vision is to create a unified cognitive-physical architecture based on LLMs and BFMs, mirroring the seamless integration of human cognition and physical control.
5.1.3 Scaling Law
@@
The concept of scaling laws describes how the performance of neural language models improves with increases in model size, data, and compute resources. While these laws are well-established in domains like language and vision, their applicability to BFMs remains an open problem [165, 166, 167, 168, 169]. Preliminary evidence suggests that scaling BFMs through larger architectures, diverse training datasets, and expanded computational resources can enhance their generalization and zero-shot adaptation capabilities. For instance, FB-CPR [46] demonstrated improved performance in humanoid control tasks with more parameters (from 25M to 288M), achieving more robust zero-shot motion tracking and reward optimization. On the other hand, data scaling appears more critical for BFMs, as it directly determines the quality and physical plausibility of the learned behavior priors, the foundational requirements that model scaling alone cannot address. However, the effect of data scaling remains underexplored, with open questions as discussed in Section 4.2.2. Furthermore, unlike LLMs, BFM scaling must strike an appropriate balance between behavior coverage (diverse motor skills) and control efficiency (precision and real-time stability). Future work should rigorously quantify these scaling dynamics to unlock BFMs’ full potential as general-purpose controllers.
5.1.4 Post-training
@@
Post-training techniques have emerged as critical tools for the continued success and refinement of foundation models, especially for LLMs [170], which refine models to improve reasoning, address limitations, and better align with user intents and ethical considerations. Among these methods, fine-tuning [171, 172, 173, 174, 175, 176, 177], integration of RL [103, 178, 179, 180, 181, 182, 183, 184], and test-time scaling [185, 186, 186, 187, 102] have been the most prominent strategies for optimizing LLMs’ performance. Integrating these post-training strategies presents unique research opportunities for BFMs. For instance, leveraging RL techniques like RLHF and RLAIF for BFMs can be crucial for refining the alignment between the agent’s behavior and real-world human expectations, especially in human-centric task environments. This opens avenues for developing more robust models that can adapt dynamically to user feedback. Additionally, test-time scaling for BFMs can optimize the computational efficiency during deployment, especially for real-time robot control or decision-making systems. Research could focus on improving the scalability of BFMs while ensuring that model outputs remain accurate and contextually appropriate across varying operational conditions.
5.1.5 Multi-agent System
@@
A multi-agent system (MAS) consists of multiple interacting agents that can be either cooperative, competitive, or a mix of both, aiming to address complex tasks that require collaboration, decision-making, and behavior coordination among agents [188, 189, 190, 191, 192, 193]. BFMs can fundamentally accelerate the construction of MAS consists of humanoid robots, eliminating the need to effortlessly teach each robot basic survival skills like balance and locomotion before they can collaborate. Instead, researchers can focus directly on higher-level coordination challenges like role allocation and team strategy. However, current BFMs trained on single-robot data lack the specialized interaction capabilities needed for optimal collaboration. This presents a promising research direction: developing next-generation BFMs trained explicitly on multi-robot interaction scenarios. Such models could better handle physical coordination challenges like object handovers, formation maintenance, and collision avoidance while preserving their generalizability.
5.1.6 Evaluation Mechanism
@@
While foundation models like LLMs benefit from well-established benchmarks (e.g., GPQA [194] for broad knowledge recall, MATH [195] for mathematical problem solving, or MUSR [196] for multi-step reasoning), there is no specific and comprehensive evaluation mechanism for evaluating BFM’s capability and guiding the evolution direction [197, 198]. A robust assessment of BFMs must consider multiple interdependent factors, including task generalization across unseen scenarios, adaptability to new skills with minimal data, robustness against physical perturbations, and alignment with human safety and interpretability standards. For example, Motivo [46] combines quantitative metrics like task success rates with qualitative human evaluations of motion naturalness, yet critical gaps remain in assessing compositional skill combinations, hardware-specific constraints, and long-term behavioral stability. Future benchmarks should focus on progressive difficulty levels and cross-domain transfer tests to effectively assess BFMs’ potential as general-purpose physical controllers in both simulated and real-world scenarios. This approach will ultimately guide the field toward developing more capable and reliable humanoid systems.
5.2 Risks
5.2.1 Ethical Issues
@@
Ethical issues consistently accompany the development of diverse foundation models [199, 200, 201], which involve biased or unlicensed data, racial discrimination, and uncontrollable behaviors, etc. For BFMs, training on non-diverse motion datasets may encode demographic biases such as favoring movements natural to specific age groups or body types, which then propagate into robotic behaviors and create embodied forms of discrimination. Meanwhile, privacy risks escalate beyond data memorization to movement analytics, where rehabilitation or performance data could leak sensitive health information through generated motions. The physical instantiation of BFMs introduces unprecedented risks: unlike purely digital models, misaligned BFMs might reproduce unsafe or socially harmful behaviors (e.g., aggressive gestures, or exclusionary motions) with real-world consequences. While techniques like differential privacy and federated learning offer partial solutions, they struggle with the continuous motion data’s temporal nature. In summary, BFMs demand novel governance frameworks that address both data provenance and normativity of real-time behavior.
5.2.2 Safety Mechanism
@@
As BFMs will be increasingly deployed in real-world robotic systems, they introduce critical safety requirements beyond those of digital foundation models [202, 203, 204]. A key issue is maintaining model behavior integrity, particularly in safety-critical scenarios such as human-robot interaction and autonomous navigation. When trained on large-scale yet weakly curated motion datasets, BFMs may unintentionally learn unsafe or undesirable behaviors. Even minor changes in sensory input—whether caused by adversarial attacks or sensor noise—can lead to control failures. This highlights the need for robustness against shifts in data distribution and protection against malicious input manipulation.
@@
While most current BFMs focus primarily on proprioceptive inputs, integrating multimodal information (e.g., visual, linguistic, and auditory cues) has emerged as a promising direction for more generalizable and situationally aware control. However, multimodality introduces new vulnerabilities. Adversaries can exploit inconsistencies across modalities, as seen in the well-known CLIP case where an apple image was misclassified as an ”iPod” due to an overlaid label [205]. Such cross-modal confusion can be especially dangerous when a BFM is adapted for an unimodal task but retains sensitivity to irrelevant signals from other modalities. These challenges underscore the need to develop robust safety mechanisms for BFMs. Future work could prioritize adversarial robustness, cross-modal consistency checks, and disentangling modality-specific information to ensure predictable, trustworthy robot behavior in open-world environments.
6 Conclusion
@@
In this paper, we present a systematic preview of the behavior foundation model, an emerging and transformative paradigm of humanoid whole-body control systems. By pre-training on large-scale and diverse humanoid behavior data, BFMs learn broad behavior priors that enable few-shot or zero-shot adaptation to diverse downstream tasks, eliminating the need for resource-intensive and task-specific retraining. We establish a comprehensive taxonomy categorizing BFM approaches into pre-training pipelines and adaptation approaches while demonstrating their real-world applicability across industrial, gaming, and healthcare domains. Furthermore, we highlight future research opportunities in multimodal BFMs, high-level ML system integration, post-training optimization, and standardized evaluation mechanisms that could accelerate BFM development.
@@
Despite their unprecedented capabilities, BFMs face significant challenges, including the Sim2Real gap, embodiment dependence, and data scarcity. The physical instantiation of BFMs introduces unique safety risks requiring robust verification mechanisms, while their training on human motion data probably raises ethical concerns regarding privacy and bias mitigation. Addressing these limitations in future work will lead to more reliable and generalizable BFMs. Our work is expected to inspire more subsequent research on BFMs.

### v4

@@ IV-B3 Data Bottleneck
The data bottleneck poses a fundamental constraint in developing BFMs for humanoid robots. While the datasets listed in Table III have been successfully employed to train the current BFMs, their scale remains significantly smaller than the datasets used to train LLMs or large vision models. This scarcity is exacerbated when retargeting motions to specific robotic platforms [gleicher1998retargetting], where subtle morphological differences can incur severe policy performance loss. Real-world robot data is even more constrained due to hardware limitations and safety concerns. The challenge compounds when considering multimodal data requirements.
@@
KIT-ML [plappert2016kit] 3911 11.2 2016
AMASS [mahmood2019amass] 11265 40.0 2019
LAFAN [harvey2020robust] 77 4.6 2020
BABEL [punnakkal2021babel] 13220 43.5 2021
Posescript [delmas2022posescript] - - 2022
HumanML3D [Guo_2022_CVPR] 14616 28.6 2022
Motion-X [lin2023motion] 81084 144.2 2023
Motion-X++ [zhang2025motion] 120462 180.9 2025
PHUMA [lee2025phuma] 76000 73.0 2025
Humanoid-X [mao2025universal] 163800 240.0 2025
@@
Current BFMs predominantly rely on proprioceptive inputs, and integrating exteroceptive sensing for real-world deployment introduces new bottlenecks. Pioneering work, such as OmniH2O [he2024omniho] explores multimodal data collection through teleoperation, pairing head-mounted RGBD cameras with whole-body proprioception and motor commands. While this enables tasks like vision-based autonomy and imitation learning, its 40-minute dataset (OmniH2O-6) remains limited to lab environments and lacks diverse contact dynamics. Currently, there is no existing dataset that provides large-scale and temporally aligned recordings of proprioception, vision, and contact dynamics across diverse environments. Therefore, larger-scale, high-quality, and more curated datasets are urgently needed to enhance the effectiveness of current BFMs and to aid in the development of future models.
IV-B4 Embodiment Generalization
@@
Despite the use of the term ”BFM”, current BFMs are typically trained on specific humanoid robot embodiments with fixed morphology, actuator dynamics, and sensor configurations. While they excel at controlling the robot they were trained on, it is currently intractable for them to generalize to novel embodiments with different body shapes, actuator types, or degrees of freedom. This limitation stems from several key challenges in embodiment generalization. One major issue is morphological mismatch, where policies designed for a specific kinematic or dynamic structure struggle to adapt to robots with differing link lengths, joint types, or mass distributions. Additionally, differences in actuator dynamics, such as variations in torque, latency, or control modes (e.g., position vs. torque control), can destabilize policies when transferred between systems. Sensor diversity introduces another layer of complexity, as BFMs often assume consistent sensor inputs and struggle to manage situations involving missing or additional sensing modalities. Furthermore, reward functions created for one type of embodiment (e.g., humanoid walking) may not effectively translate to others (e.g., hexapod gait patterns), resulting in confusion during task adaptation. These challenges underscore the necessity for more adaptable BFM architectures that can abstract skills across various robotic platforms.
V Opportunities and Risks
@@
In this section, we explore key research opportunities that could enhance the broader adoption of BFMs, including their integration into high-level ML systems, scaling laws, post-training techniques, and multi-agent coordination. We also critically examine the risks associated with BFMs, such as ethical concerns in human-robot interaction and safety mechanisms for real-world deployment. An overview of the opportunities and risks is illustrated in Fig. 10. By addressing these opportunities and risks, the robotics community can ensure that BFMs evolve into robust, adaptable, and socially responsible controllers for next-generation humanoid systems.
@@
V-A Opportunities
V-A1 Multimodal BFMs
@@
A promising direction for BFMs lies in expanding beyond proprioceptive inputs to multimodal sensory integration, incorporating exteroceptive signals such as vision, acoustic signals, and tactile feedback [liu2022multimodal, yao2024multimodal, wang2024multimodal]. While current BFMs primarily rely on proprioceptive information, richer perceptual inputs could enable more robust and adaptive behaviors in unstructured environments. For example, integrating real-time visual perception could enable humanoids to dynamically adjust their movements based on object positions, terrain conditions, or human interactions, enhancing both safety and task performance. However, achieving multimodal BFMs requires the advancement of both large-scale datasets and scalable training paradigms, which are key challenges for the field.
V-A2 High-level ML System
@@
Recent advances in foundation models like LLMs and VLAs have demonstrated their potential as controllers for coordinating specialized AI systems [shen2023hugginggpt, dai2023llmintheloop, chan2024chateval, maeureka, liu2025aligning, yu2025prophet, du2025reinforcement]. For example, HuggingGPT [shen2023hugginggpt] leverages LLMs to orchestrate task planning across diverse models, while Eureka [maeureka] employs LLMs to automate reward function design in RL. These projects present a natural pathway for integrating BFMs into high-level ML systems, where they could serve as universal low-level controllers for humanoid robots. By combining the reasoning capabilities of LLMs (for task planning and adaptation) with the motor priors encoded in BFMs (for real-time execution), such systems could achieve unprecedented flexibility in handling complex and multi-step physical tasks, while minimizing task-specific engineering. The ultimate vision is to create a unified cognitive-physical architecture based on LLMs and BFMs, mirroring the seamless integration of human cognition and physical control.
V-A3 Scaling Law
@@
The concept of scaling laws describes how the performance of neural language models improves with increases in model size, data, and compute resources. While these laws are well-established in domains like language and vision, their applicability to BFMs remains an open problem [kaplan2020scaling, clark2022unified, aghajanyan2023scaling, isik2024scaling, que2024d]. Preliminary evidence suggests that scaling BFMs through larger architectures, diverse training datasets, and expanded computational resources can enhance their generalization and zero-shot adaptation capabilities. For instance, FB-CPR [tirinzoni2025zeroshot] demonstrated improved performance in humanoid control tasks with more parameters (from 25M to 288M), achieving more robust zero-shot motion tracking and reward optimization. On the other hand, data scaling appears more critical for BFMs, as it directly determines the quality and physical plausibility of the learned behavior priors, the foundational requirements that model scaling alone cannot address. However, the effect of data scaling remains underexplored, with open questions as discussed in Section IV-B3. Furthermore, unlike LLMs, BFM scaling must strike an appropriate balance between behavior coverage (diverse motor skills) and control efficiency (precision and real-time stability). Future work should rigorously quantify these scaling dynamics to unlock the full potential of BFMs as general-purpose controllers.
V-A4 Post-training
@@
Post-training techniques have emerged as critical tools for the continued success and refinement of foundation models, especially for LLMs [kumar2025llm], which refine models to improve reasoning, address limitations, and better align with user intents and ethical considerations. Among these methods, fine-tuning [yue2023disc, chenlonglora, luong2024reft, ding2023parameter, malladi2023fine, xu2023parameter, guo2024tuning], integration of RL [ouyang2022training, yuan2023rrhf, liu2024skywork, dong2024rlhf, cui2023ultrafeedback, du2023guiding, song2023self, xu2025magpie], and test-time scaling [jiang2014searching, yao2023tree, yao2023tree, tian2024toward, gandhi2024stream] have been the most prominent strategies for optimizing LLMs’ performance. Integrating these post-training strategies presents unique research opportunities for BFMs. For instance, leveraging RL techniques like RLHF and RLAIF for BFMs can be crucial for refining the alignment between the agent’s behavior and real-world human expectations, especially in human-centric task environments. This opens avenues for developing more robust models that can adapt dynamically to user feedback. Additionally, test-time scaling for BFMs can enhance computational efficiency during deployment, particularly for real-time robot control or decision-making systems. Research could focus on improving the scalability of BFMs while ensuring that model outputs remain accurate and contextually appropriate across varying operational conditions.
V-A5 Multi-agent System
@@
A multi-agent system (MAS) consists of multiple interacting agents that can be either cooperative, competitive, or a mix of both, aiming to address complex tasks that require collaboration, decision-making, and behavior coordination among agents [arai2002advances, gautam2012review, dorri2018multi, madridano2021trajectory, zhou2021survey, ju2022review]. BFMs can fundamentally accelerate the construction of MAS consists of humanoid robots, eliminating the need to effortlessly teach each robot basic survival skills like balance and locomotion before they can collaborate. Instead, researchers can focus directly on higher-level coordination challenges, such as role allocation and team strategy. However, current BFMs trained on single-robot data lack the specialized interaction capabilities needed for optimal collaboration. This presents a promising research direction: developing next-generation BFMs trained explicitly on multi-robot interaction scenarios. Such models could better handle physical coordination challenges, such as object handovers, formation maintenance, and collision avoidance, while preserving their generalizability.
V-A6 Evaluation Mechanism
@@
While foundation models like LLMs benefit from well-established benchmarks (e.g., GPQA [rein2024gpqa] for broad knowledge recall, MATH [hendrycks2measuring] for mathematical problem solving, or MUSR [spraguemusr] for multi-step reasoning), there is no specific and comprehensive evaluation mechanism for evaluating BFM’s capability and guiding the evolution direction [xu2024lvlm, huang2025causality]. A robust assessment of BFMs must consider multiple interdependent factors, including task generalization across unseen scenarios, adaptability to new skills with minimal data, robustness against physical perturbations, and alignment with human safety and interpretability standards. For example, Motivo [tirinzoni2025zeroshot] combines quantitative metrics, such as task success rates, with qualitative human evaluations of motion naturalness; yet, critical gaps remain in assessing compositional skill combinations, hardware-specific constraints, and long-term behavioral stability. Future benchmarks should focus on progressive difficulty levels and cross-domain transfer tests to effectively assess the potential of BFMs as general-purpose physical controllers in both simulated and real-world scenarios. This approach will ultimately guide the field toward developing more capable and reliable humanoid systems.
V-B Risks
V-B1 Ethical Issues
@@
Ethical issues consistently accompany the development of diverse foundation models [weidinger2021ethical, yan2024practical, bommasani2024considerations], which involve biased or unlicensed data, racial discrimination, and uncontrollable behaviors, etc. For BFMs, training on non-diverse motion datasets may encode demographic biases such as favoring movements natural to specific age groups or body types, which then propagate into robotic behaviors and create embodied forms of discrimination. Meanwhile, privacy risks escalate beyond data memorization to movement analytics, where rehabilitation or performance data could leak sensitive health information through generated motions. The physical instantiation of BFMs introduces unprecedented risks: unlike purely digital models, misaligned BFMs might reproduce unsafe or socially harmful behaviors (e.g., aggressive gestures, or exclusionary motions) with real-world consequences. While techniques like differential privacy and federated learning offer partial solutions, they struggle with the temporal nature of continuous motion data. In summary, BFMs demand novel governance frameworks that address both data provenance and normativity of real-time behavior.
V-B2 Safety Mechanism
@@
As BFMs will be increasingly deployed in real-world robotic systems, they introduce critical safety requirements beyond those of digital foundation models [vasic2013safety, lasota2017survey, kumagai2019toward]. A key issue is maintaining model behavior integrity, particularly in safety-critical scenarios such as human-robot interaction and autonomous navigation. When trained on large-scale yet weakly curated motion datasets, BFMs may unintentionally learn unsafe or undesirable behaviors. Even minor changes in sensory input, caused by adversarial attacks or sensor noise, can lead to control failures. This highlights the need for robustness against shifts in data distribution and protection against malicious input manipulation.
@@
While most current BFMs focus primarily on proprioceptive inputs, integrating multimodal information (e.g., visual, linguistic, and auditory cues) has emerged as a promising direction for more generalizable and situationally aware control. However, multimodality introduces new vulnerabilities. Adversaries can exploit inconsistencies across modalities, as seen in the well-known CLIP case where an apple image was misclassified as an ”iPod” due to an overlaid label [goh2021multimodal]. Such cross-modal confusion can be especially dangerous when a BFM is adapted for an unimodal task but retains sensitivity to irrelevant signals from other modalities. These challenges underscore the need to develop robust safety mechanisms for BFMs. Future work could prioritize adversarial robustness, cross-modal consistency checks, and disentangling modality-specific information to ensure predictable, trustworthy robot behavior in open-world environments.
VI Conclusion
@@
In this paper, we present a systematic review of the behavior foundation model, an emerging and transformative paradigm of humanoid whole-body control systems. By pre-training on large-scale and diverse humanoid behavior data, BFMs learn broad behavior priors that enable few-shot or zero-shot adaptation to diverse downstream tasks, eliminating the need for resource-intensive and task-specific retraining. We establish a comprehensive taxonomy that categorizes BFM approaches into pre-training pipelines and adaptation strategies, demonstrating their real-world applicability across industrial, gaming, and healthcare domains. Furthermore, we highlight future research opportunities in multimodal BFMs, high-level ML system integration, post-training optimization, and standardized evaluation mechanisms that could accelerate BFM development.
@@
Despite their unprecedented capabilities, BFMs face significant challenges, including the Sim2Real gap, embodiment dependence, and data scarcity. The physical instantiation of BFMs introduces unique safety risks that require robust verification mechanisms, while their training on human motion data likely raises ethical concerns regarding privacy and bias mitigation. Addressing these limitations in future work will lead to more reliable and generalizable BFMs. Our work is expected to inspire more subsequent research on BFMs.